# NeuSDFusion: A Spatial-Aware Generative Model for 3D Shape Completion, Reconstruction, and Generation

### v1

@@ Abstract
3D shape generation aims to produce innovative 3D content adhering to specific conditions and constraints. Existing methods often decompose 3D shapes into a sequence of localized components, treating each element in isolation without considering spatial consistency. As a result, these approaches exhibit limited versatility in 3D data representation and shape generation, hindering their ability to generate highly diverse 3D shapes that comply with the specified constraints. In this paper, we introduce a novel spatial-aware 3D shape generation framework that leverages 2D plane representations for enhanced 3D shape modeling. To ensure spatial coherence and reduce memory usage, we incorporate a hybrid shape representation technique that directly learns a continuous signed distance field representation of the 3D shape using orthogonal 2D planes. Additionally, we meticulously enforce spatial correspondences across distinct planes using a transformer-based autoencoder structure, promoting the preservation of spatial relationships in the generated 3D shapes. This yields an algorithm that consistently outperforms state-of-the-art 3D shape generation methods on various tasks, including unconditional shape generation, multi-modal shape completion, single-view reconstruction, and text-to-shape synthesis.
@@ 1 Introduction
Recent studies [18, 13, 35] have focused on understanding 3D geometry by leveraging multi-view supervision with predefined camera positions and incorporating inductive biases through neural rendering [30] techniques. While these approaches yield impressive results, the training process can be time-consuming and often overlooks accessible 3D data that could be harnessed to derive more effective shape priors. To directly exploit such shape priors from 3D data, contemporary methods [7, 23] typically model the distribution of 3D objects using explicit 3D representations. These techniques frequently employ Truncated Signed Distance Function (T-SDF) as the 3D object representation and utilize an encoder to embed an object directly into a latent representation. However, the T-SDF representation is generally memory-intensive, hindering the capture of detailed shape characteristics due to memory constraints. Furthermore, T-SDF truncates distance values, leading to a discontinuous representation that may not be optimal for applications requiring smooth surfaces or continuous fields. Inspired by recent efforts [4] in using 2D plane representations for 3D shape modeling, some works [9, 15] opt to encode an object as a tri-plane representation for embedding and employ point clouds as an intermediate representation. However, such an intermediate representation necessitates extra encoding networks, resulting in reduced efficiency. NFD [44], a work closely related to ours, utilizes tri-planes as the 3D object representation. However, their tri-planes are learned from occupancy grids, so that the representation capability is constrained by the grid resolution. This limitation highlights the need for more advanced 3D shape representations that can effectively capture spatial coherence and generate high-quality 3D shapes, paving the way for further progress in the field of 3D generative models.
@@
To this end, we present NeuSDFusion, a framework for generating high-fidelity 3D objects, as depicted in Fig. 1. Unlike previous methods that either rely on rendering-based techniques, ignoring available explicit 3D priors, or employ raw 3D data (point clouds [49], mesh [25], occupancy grid [44], or SDF [31, 23]) as 3D representations, limited by memory constraints, we introduce a hybrid tri-plane SDF representation named NeuSDF. This representation expressively embeds 3D objects as orthogonal 2D planes, preserving 3D structural information and generating smooth surfaces, while also being computationally efficient and capable of handling high-resolution details.
@@ 2.1 Generative Modeling of 3D Shapes
In recent years, numerous works have been proposed to generate 3D shapes, with existing 3D generative models built on various frameworks. This includes generative adversarial networks (GANs)[52, 1, 60], variational autoencoders (VAEs)[48, 14, 21, 31], normalizing flows [58, 31], autoregressive models [47, 31], energy-based models [56, 11], and more recently, denoising diffusion probabilistic models (DDPMs) [29, 62, 25, 49, 23, 7, 44, 54]. Luo et al. [29] pioneered the application of DDPMs for modeling raw point clouds. Building on the success of latent diffusion models [42] in 2D image generation, several works [49, 23, 7] explored generative modeling of 3D shapes in latent space to reduce computational complexity and enhance generation quality. Within this context, LION [49] utilized point clouds to represent 3D objects, while 3DQD [23], SDFusion [7], and DiffusionSDF [9] employed the Signed Distance Function (SDF) for 3D shape representation. Concurrently, other research efforts have investigated DDPMs for 3D shape generation with varied representations, such as mesh [25], occupancy grid [44], and neural radiance fields [32]. Despite these advancements, a common limitation among most approaches is their inadvertent neglect of spatial coherence during the generation process. Addressing this gap, a recent study by Wang et al. [51] introduces a method that maintains spatial correspondence through the use of 3D-aware convolution on rolled-out tri-planes. Nevertheless, this technique employs axis-wise pooling to downsample plane features to a vector and only aggregates such downsampled features without enforcing in-plane communication. As a result, it lacks context information and fails to generate smooth shape surfaces. This highlights the need for more advanced methods that can effectively capture spatial coherence and generate high-quality 3D shapes.
@@ 2.2 Representation of 3D Shapes
Various 3D shape representations have been investigated to facilitate the 3D synthesis task. We categorize them into two primary classes: rendering-based and rendering-free methods. Rendering-based approaches, such as GET3D [13], Zero123 [24], employ multi-view images to learn object geometry via volume rendering [30]. However, these methods overlook the available 3D objects as priors, leading to coarse shapes that often lack fine geometry details. On the other hand, rendering-free methods model the distribution of 3D objects using raw 3D representations, such as point clouds, mesh, and T-SDF. These methods necessitate careful network design (e.g., treating point clouds as sets [29] or handling mesh edges [25]), and may struggle to represent intricate object structures. SDFusion [7] and 3DQD [23] employ T-SDF to represent 3D objects. However, direct use 3D representation requires substantial memory, while most of the 3D space remains devoid of shape surfaces. Consequently, these 3D representation approaches are generally less effective compared to 2D plane formulations [4]. Moreover, T-SDF truncates distance values, leading to a discontinuous representation more sensitive to noise compared to SDF, and not ideal for applications requiring smooth surfaces or continuous fields, which are often primary requirements for 3D shape generation. NDF [44] was the first to introduce a tri-plane representation for modeling raw 3D data. However, their tri-plane representation was derived from a discrete occupancy grid, constraining shape details by the resolution of planes.
@@ 3 Methodology
Figure 2: Our method follows a pipeline consisting of three stages. Initially, given a raw mesh, we sample surface points and space-filling points in the first stage to adapt each mesh to a NeuSDF tri-plane representation. In the second stage, we compress the raw tri-plane representation into latent tri-planes $z$ by employing a spatial-aware autoencoder. In the third stage, we train a latent diffusion model capable of generating tri-plane latent $z_{0}$ from a standard Gaussian under flexible conditions. During the inference phase, we input the generated latent $z_{0}$ into the decoder, and generate a mesh using the Marching Cubes algorithm by querying the signed distance value of any position via interpolating the reconstructed tri-plane.
Our proposed pipeline comprises three stages, as illustrated in Fig. 2. Initially, we model each 3D object with a novel hybrid representation named NeuSDF, mapping the detailed signed distance function into three orthogonal 2D planes (Section 3.1). We demonstrate that NeuSDF delivers superior 3D synthesis quality with improved computational efficiency. In the second stage, we compress this 2D plane representation into latent features. To this end, we propose a transformer-based autoencoder, which efficiently encodes the tri-planes into a compact representation while preserving the 3D correlations between planes (Section 3.2). In the final stage, we train a diffusion model capable of performing unconditional or guided generation with prompts from various modalities, including images, text, and point clouds (Section 3.3).
@@ 3.1 Representing objects as NeuSDF
Formulation. A signed distance function is a function that maps a spatial coordinate $(x,y,z)$ to a real value $s$, indicating the distance of the spatial position to the object surface. This function is commonly represented with a network or a vector representation [33, 31]. Instead, inspired by the recent success of the tri-plane representation [4], we represent the geometry of a 3D shape using three axis-aligned planes, i.e. the XY, YZ, and XZ planes. We then use a multi-layer perceptron (MLP) to decode the tri-plane into signed distance values. Formally, we query the signed distance value of any 3D position $p\in\mathbb{R}^{3}$ by projecting it onto each of the three feature planes, retrieving the corresponding feature vector $F_{xy},F_{xz},F_{yz}$ via bilinear interpolation and aggregating the three feature vectors to obtain the feature vector for the query position. Then, we use a MLP $\phi$ to interpret the aggregated feature as a real value. The signed distance of any query position $p$ is acquired as follows:
@@
Point Sampling. To effectively represent a 3D object as a NeuSDF, we sample SDF values to train the tri-planes. Given a 3D object mesh, we first preprocess it to ensure it forms a watertight shape using the method proposed by Huang et al. [19]. Then, we normalize the shape within the $[-1,1]^{3}$ box. Subsequently, we randomly sample on-surface points $\Omega_{0}$ from the shape surface and uniformly sample off-surface points $\Omega$ with ground truth SDF values in the $[-1,1]^{3}$ space. In addition to the signed distance supervision, we leverage normal direction as extra guidance to achieve detailed surface modeling. Consequently, we also sample the normal vector for each on-surface point.
@@
where $d_{p}$ and $n_{p}$ are ground truth SDF value and surface normal vector, respectively. The gradient $\nabla_{p}\Phi(p)=\left(\frac{\partial\Phi(p)}{\partial X},\frac{\partial\Phi(p)}{\partial Y},\frac{\partial\Phi(p)}{\partial Z}\right)$ represents the direction of the steepest change in SDF. It can be computed using finite difference, e.g., the partial derivative for the X-axis component reads $\frac{\partial\Phi(p)}{\partial X}=\frac{\Phi(p+(\delta,0,0))-\Phi(p-(\delta,0,0))}{2\delta}$ where $\delta$ is the step size. The Eikonal loss constrains $\|\nabla_{p}\Phi(p)\|$ to be $1$ almost everywhere, thus maintaining the intrinsic physical property of the signed distance function.
@@ 3.3 Generative Modeling of Latent Tri-Planes
We employ the denoising U-Net architecture described in Stable Diffusion [42], augmenting it with an additional cross-attention layer within each block of the U-Net backbone to incorporate condition prompts. The cross-attention layer is defined as: $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$, with:
@@ 4.4 Language-Guided Generation
Figure 6: Text-guided shape generation. We showcase the ability of our method w.r.t. generating shapes based on text prompts. The orange text represents the geometry prompt and we employ Meshy [^2] https://www.meshy.ai/to generate texture using the blue prompt.
@@ 4.5 Ablation Study
We evaluate the effectiveness of our spatial-aware transformer-based autoencoder design by comparing it with two related methods: roll out [15, 51] and channel-wise concatenation [44, 9]. Specifically, we train our autoencoder and a CNN-based autoencoder from Stable Diffusion [42] using these methods on a chair-class subset of ShapeNet. As illustrated in Fig. 7, the reconstructed results of channel-wise concatenation or roll out exhibit defects due to spatial misalignment. The roll out method suffers from defects at the border of the tri-plane representation, as the convolution operation overlaps two planes,
@@
while the values at the edge of two adjacent planes are not continuous in space, as shown in Fig. 7 (b). Channel concatenation, in contrast, presents some reconstruction defects due to the lack of an explicit relationship between the same spatial locations across different planes, thereby limiting the reconstruction capability. Our proposed approach employs group convolution to downsample each plane individually and introduces a specially designed attention mechanism to achieve 3D-aware tri-plane interaction. Consequently, our generated objects are integral with a smooth surface. Furthermore, we provide additional evaluation and implementation details of our method in the supplementary material.

### v2

@@ Abstract
3D shape generation aims to produce innovative 3D content adhering to specific conditions and constraints. Existing methods often decompose 3D shapes into a sequence of localized components, treating each element in isolation without considering spatial consistency. As a result, these approaches exhibit limited versatility in 3D data representation and shape generation, hindering their ability to generate highly diverse 3D shapes that comply with the specified constraints. In this paper, we introduce a novel spatial-aware 3D shape generation framework that leverages 2D plane representations for enhanced 3D shape modeling. To ensure spatial coherence and reduce memory usage, we incorporate a hybrid shape representation technique that directly learns a continuous signed distance field representation of the 3D shape using orthogonal 2D planes. Additionally, we meticulously enforce spatial correspondences across distinct planes using a transformer-based autoencoder structure, promoting the preservation of spatial relationships in the generated 3D shapes. This yields an algorithm that consistently outperforms state-of-the-art 3D shape generation methods on various tasks, including unconditional shape generation, multi-modal shape completion, single-view reconstruction, and text-to-shape synthesis. Our project page is available at https://weizheliu.github.io/NeuSDFusion/.
@@ 1 Introduction
Recent studies [19, 15, 36] have focused on understanding 3D geometry by leveraging multi-view supervision with predefined camera positions and incorporating inductive biases through neural rendering [31] techniques. While these approaches yield impressive results, the training process can be time-consuming and often overlooks accessible 3D data that could be harnessed to derive more effective shape priors. To directly exploit such shape priors from 3D data, contemporary methods [8, 24] typically model the distribution of 3D objects using explicit 3D representations. These techniques frequently employ Truncated Signed Distance Function (T-SDF) as the 3D object representation and utilize an encoder to embed an object directly into a latent representation. However, the T-SDF representation usually employs a 3D voxel representation to store the distance field, which is generally memory-intensive, hindering the capture of detailed shape characteristics due to memory constraints. Inspired by recent efforts [4] in using 2D plane representations for 3D shape modeling, some works [9, 17] opt to encode an object as a tri-plane representation for embedding and employ point clouds as an intermediate representation. However, such an intermediate representation necessitates extra encoding networks, resulting in reduced efficiency. NFD [43], a work closely related to ours, utilizes tri-planes as the 3D object representation. However, their tri-planes are learned from occupancy grids, so that the representation capability is constrained by the grid resolution. This limitation highlights the need for more advanced 3D shape representations that can effectively capture spatial coherence and generate high-quality 3D shapes, paving the way for further progress in the field of 3D generative models.
@@
To this end, we present NeuSDFusion, a framework for generating high-fidelity 3D objects, as depicted in Fig. 1. Unlike previous methods that either rely on rendering-based techniques, ignoring available explicit 3D priors, or employ raw 3D data (point clouds [46], mesh [26], occupancy grid [43, 35], or SDF [32, 24]) as 3D representations, limited by memory constraints, we introduce a hybrid tri-plane SDF representation named NeuSDF. This representation expressively embeds 3D objects as orthogonal 2D planes, preserving 3D structural information and generating smooth surfaces, while also being computationally efficient and capable of handling high-resolution details.
@@ 2.1 Generative Modeling of 3D Shapes
In recent years, numerous works have been proposed to generate 3D shapes, with existing 3D generative models built on various frameworks. This includes generative adversarial networks (GANs)[48, 1, 61], variational autoencoders (VAEs)[45, 16, 22, 32], normalizing flows [54], autoregressive models [57], energy-based models [51, 11], and more recently, denoising diffusion probabilistic models (DDPMs) [63, 49, 6, 13, 20]. Luo et al. [30] pioneered the application of DDPMs for modeling raw point clouds. Building on the success of latent diffusion models [41] in 2D image generation, several works [46, 24, 8] explored generative modeling of 3D shapes in latent space to reduce computational complexity and enhance generation quality. Within this context, LION [46] utilized point clouds to represent 3D objects, while 3DQD [24], SDFusion [8], and DiffusionSDF [9] employed the Signed Distance Function (SDF) for 3D shape representation. Concurrently, other research efforts have investigated DDPMs for 3D shape generation with varied representations, such as mesh [26], occupancy grid [43], and neural radiance fields [33]. Despite these advancements, a common limitation among most approaches is their inadvertent neglect of spatial coherence during the generation process. Addressing this gap, a recent study by Wang et al. [47] introduces a method that maintains spatial correspondence through the use of 3D-aware convolution on rolled-out tri-planes. Nevertheless, this technique employs axis-wise pooling to downsample plane features to a vector and only aggregates such downsampled features without enforcing in-plane communication. As a result, it lacks context information and fails to generate smooth shape surfaces. This highlights the need for more advanced methods that can effectively capture spatial coherence and generate high-quality 3D shapes.
@@ 2.2 Representation of 3D Shapes
Various 3D shape representations have been investigated to facilitate the 3D synthesis task. We categorize them into two primary classes: rendering-based and rendering-free methods. Rendering-based approaches [21, 19], such as LRM [19], Zero123 [25], employ multi-view images to learn object geometry via volume rendering [31]. However, these methods overlook the available 3D objects as priors, leading to coarse shapes that often lack fine geometry details. On the other hand, rendering-free methods model the distribution of 3D objects using raw 3D representations, such as point clouds [12], mesh [3], binary occupancy [53, 58], and raw SDF [60, 59, 55]. These methods necessitate careful network design (e.g., treating point clouds as sets [30] or handling mesh edges [26]), and may struggle to represent intricate object structures. SDFusion [8] and 3DQD [24] employ T-SDF to represent 3D objects. However, direct use 3D representation requires substantial memory, while most of the 3D space remains devoid of shape surfaces. Consequently, these 3D representation approaches are generally less effective compared to 2D plane formulations [4]. NDF [43] was the first to introduce a tri-plane representation for modeling raw 3D data. However, their tri-plane representation was derived from a discrete occupancy grid, constraining shape details by the resolution of planes.
@@ 3 Methodology
Figure 2: Our method follows a pipeline consisting of three stages. Given a raw mesh, we first sample surface points and space-filling points to adapt each mesh to a NeuSDF representation. In the second stage, we compress the raw tri-plane representation into latent tri-planes $z$ with a spatial-aware autoencoder. In the third stage, we train a latent diffusion model capable of generating tri-plane latent $z_{0}$ from a standard Gaussian under flexible conditions. During the inference phase, we input the generated latent $z_{0}$ into the decoder, and generate a mesh using the Marching Cubes algorithm by querying the signed distance value of any position via interpolating the reconstructed tri-plane.
Our pipeline comprises three stages, as illustrated in Fig. 2. Initially, we model each 3D object with a novel hybrid representation named NeuSDF, mapping the signed distance function into three orthogonal 2D planes (Sec. 3.1).
In the second stage, we compress this 2D plane representation into latent features. To this end, we propose a transformer-based autoencoder, which efficiently encodes the tri-planes into a compact representation while preserving the 3D correlations between planes (Sec. 3.2). In the final stage, we train a diffusion model capable of performing unconditional or guided generation with prompts from various modalities, including images, text, and point clouds (Sec. 3.3).
@@ 3.1 Representing objects as NeuSDF
Formulation. A signed distance function is a function that maps a spatial coordinate $(x,y,z)$ to a real value $s$, indicating the distance of the spatial position to the object surface. This function is commonly represented with a network or a vector representation [34, 32]. Instead, inspired by the recent success of the tri-plane representation [4], we represent the geometry of a 3D shape using three axis-aligned planes, i.e. the XY, YZ, and XZ planes. We then use a multi-layer perceptron (MLP) to decode the tri-plane into signed distance values. Formally, we query the signed distance value of any 3D position $p\in\mathbb{R}^{3}$ by projecting it onto each of the three feature planes, retrieving the corresponding feature vector $F_{xy},F_{xz},F_{yz}$ via bilinear interpolation and aggregating the three feature vectors to obtain the feature vector for the query position. Then, we use an MLP $\phi$ to interpret the aggregated feature as a real value. The signed distance of any query position $p$ is acquired as follows:
@@
Point Sampling. To effectively represent a 3D object as a NeuSDF, we sample SDF values to train the tri-planes. Given a 3D object mesh, we first preprocess it to transform it to a watertight shape using Blender’s voxel remeshing tool. Then, we normalize the shape within the $[-1,1]^{3}$ box. Subsequently, we randomly sample on-surface points $\Omega_{0}$ from the shape surface and uniformly sample off-surface points $\Omega$ with ground truth SDF values in the $[-1,1]^{3}$ space. In addition to the signed distance supervision, we leverage normal direction as extra guidance to achieve detailed surface modeling. Consequently, we also sample the normal vector for each on-surface point.
@@
where $d_{p}$ and $n_{p}$ are ground truth SDF value and surface normal vector, respectively. The gradient $\nabla_{p}\Phi(p)=\left(\frac{\partial\Phi(p)}{\partial X},\frac{\partial\Phi(p)}{\partial Y},\frac{\partial\Phi(p)}{\partial Z}\right)$ represents the direction of the steepest change in SDF. It can be computed using finite differences, e.g., the partial derivative for the X-axis component reads $\frac{\partial\Phi(p)}{\partial X}=\frac{\Phi(p+(\delta,0,0))-\Phi(p-(\delta,0,0))}{2\delta}$ where $\delta$ is the step size. The Eikonal loss constrains $\|\nabla_{p}\Phi(p)\|$ to be $1$ almost everywhere, thus maintaining the intrinsic physical property of the signed distance function.
@@ 3.3 Generative Modeling of Latent Tri-Planes
We employ the denoising U-Net architecture described in Rombach et al. [41], augmenting it with an additional cross-attention layer within each block of the U-Net backbone to incorporate condition prompts. The cross-attention layer is defined as: $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$, with:
@@ 4.4 Language-Guided Generation
Figure 6: Text-guided shape generation. We showcase the ability of our method w.r.t. generating shapes based on text prompts. The orange text represents the geometry prompt and we employ Meshyhttps://www.meshy.ai/to generate texture using the blue prompt.
@@ 4.5 Ablation Study
We evaluate the effectiveness of our spatial-aware transformer-based autoencoder design by comparing it with two related methods: roll out [17, 47] and channel-wise concatenation [43, 9]. Specifically, we train both our autoencoder and a CNN-based autoencoder from Stable Diffusion [41] on a chair-class subset of ShapeNet. As illustrated in Fig. 7, the reconstructed results of channel-wise concatenation or roll out exhibit defects due to spatial misalignment.
@@
The roll out method suffers from defects at the border of the tri-plane representation, as the convolution operation overlaps over two planes, while the values at the edge of two adjacent planes are not continuous in space, as shown in Fig. 7 (b). Channel concatenation, in contrast, presents some reconstruction defects due to the lack of an explicit relationship between the same spatial locations across different planes, thereby limiting the reconstruction capability. Our proposed approach employs group convolution to downsample each plane individually and introduces a specially designed attention mechanism to achieve 3D-aware tri-plane interaction. Consequently, our generated objects are integral with a smooth surface. Furthermore, we provide additional evaluation and implementation details of our method in the supplementary material.